OpenStudy (anonymous):

175% as a fraction and decimal

OpenStudy (karatechopper):

1 3/4 ans 1.75

OpenStudy (mertsj):

percent means per hundred so 175% means 175/100=7/4=1 3/4 = 1.75

OpenStudy (karatechopper):

oops i mean and

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (karatechopper):

i gots it right yay!!! and can u answer on chat?